OpenStudy (anonymous):

Use Cramer's Rule to solve the system or to determine that the system is inconsistent or contains dependedent equations. 3x+ 2y=11 and 2x-3y=3 1) dependent equations 2) {(32, 8)} 3) inconsistent system 4) {(28, 32)}

OpenStudy (anonymous):

\[3x+2y=11\] \[2x-3y=3\] cramer's rule i guess uses the determinant. it is not zero because it is \[3\times (-3)-2\times 2=-9-4=-13\]

OpenStudy (anonymous):

however the solution is \(x=3,y=1\) so there is a typo somewhere

OpenStudy (anonymous):

smh thats interesting cause this was on one of my test but thanks for your help . @satellite73

OpenStudy (anonymous):

yw